SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Nov.22, 2005)

Attendees: Rainer Bartoldus, Prafulla Behera, Kris Chaisanguanthum, Selina Li, Dong Su, Dan Walker.

Operations

  • PEP is scrubbing the beam pipe and BaBar is taking data only when background gets reasonable. There were a few physics low lumi runs taken over the past week weekend and Monday. We were asked to check and watch for data quality of these runs.

EMT

  • we would need a PR release in before changing the EMT prescale. It makes sense to request a cold build. Dan is working on getting a histogram in L1TFMon to check if the prescale is really working (e.g. histogram prescaled and non-prescaled events in separate bins). Dan will give Jim Hamilton a tag for fastmon.
  • Dan will check the turn-on curve again.

DCT/DCZ

  • Marc had made some progress in getting a new version of the firmware to test last week. It had problem recovering from fiber being pulled (failed 20% of the tests). This is still under investigation.

L3

  • Prafulla fixed the L3 efficiency plots for OPR. See plots shown.
  • Rainer worked on getting the trigger mask move to trgconfig. Every run type now has a trgconfig associate with it, including the calibration runs.
  • Rainer also worked on the problem with the TSF timing histogram which had conflict with offline check. This was tested and he will commit it either today or tomorrow to get it in the next release. Rainer also fixed a few bugs here and there along the way.

Data Quality and Monitoring

  • There wasn't a definite conclusion on the runs discussed last week for the reprocessed runs from run4.
  • The low lumi runs from run5b were looked at. Overall they looked okay from the fastmon plots. However, the plots Selina showed in the meeting indicated low efficiencies in superlayer 1,4,7,and 10 when compared to runs taken in run5a. Su Dong and Rainer suggested to look at hadronic events as well as EMT triggered events. More plots were added after the meeting to look at these with some event selections. Su Dong pointed out that it looked like a DCH problem since the axial layers are in superlayer 1,4,7,10 and TSF doesn't know about axial layers. See HN posted after the operation meeting.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.